# S-Graphs 2.0 – A Hierarchical-Semantic Optimization and Loop Closure for SLAM

### v2

@@ VI-A Methodology
VI-B Results and Discussions
Single-Floor Dataset. The results of the single-floor experiments, in Table. I, show that our S-Graphs 2.0 is able to provide final map accuracy similar to the second best baseline S-Graphs+ with an average accuracy improvement of . However, the benefits of the hierarchical optimization of our S-Graphs 2.0 clearly stand out in Table. IV, where it has an average cost reduction of to the closest baseline. The benefits of the hierarchical optimization become even more clear for longer sequences. See, for example, how in C2F2, C3F2 there is slight decrease in the accuracy with respect to S-Graphs+ ( and respectively),but a large reduction the computation time ( and respectively). Summing up, in these single-floor datasets, we show that the optimization strategy in S-Graphs 2.0 maintains the estimation accuracy while significantly reducing the computational cost.
Computation Time | |||
Dataset | Seq. Length [s] | S-Graphs+ [ms] | Ours [ms] |
@@ VI-B Results and Discussions
Multi-Floor Dataset. Table. II shows the Mean Map Entropy (MME) for the multi-floor dataset. Our method outperforms all the baselines by an average of when comparing the MME. Given the large area covered in these multi-floor experiments, all baselines fail at giving reasonable estimates, being unable to maintain the separation between the floor levels as well as accumulating significant drift. This can be assessed in Fig. 8, that shows the maps estimated by S-Graphs 2.0 and S-Graphs+ for LC1. As this environment is highly aliased between floors, S-Graphs+’s loop closure gives a high number of false positives between levels. Our approach, given the floor segmentation, stairway detection, and floor-based loop closure is able to maintain the accurate map while segmenting the keyframes and its connected layers based on the given floor-level. Fig. 9 shows qualitative results for dataset CS2 demonstrating that our method is capable of segmenting the three floor-levels while maintaining a good map accuracy compared to S-Graphs+ (which again performs inaccurate loop closures between floor levels) and BALM. Table. III and Fig. 7 also validate the proper floor segmentation by comparing the IoU percentage of (height) values per floor-level from the maps generated by the algorithms versus the percentage ground truth time spend per floor-level. Observe that the IoU of our S-Graphs 2.0 correlates perfectly with the ground truth time, BALM presents a weaker correlation and the rest of baselines offer very poor results.
Furthermore, as can be seen from Table. IV the average reduction of computation time in multi-floor dataset of our approach compared to S-Graphs+ is an average of given the fact that S-Graphs+ does not perform efficient computation with increased length of the sequence as compared to our approach which efficiently handles and optimizes the keyframes and its connected semantic-relations based on the local, floor-local and room-level local optimizations while performing accurate floor-based loop closures. Additional qualitative results: https://www.youtube.com/watch?v=vOH8X7XnQms.
@@ VII Conclusion
In this work, we present S-Graphs 2.0, a multi-layered hierarchical SLAM for large, multi-floor indoor environments. Our approach leverages semantic relationships and a robust floor detection module to assign floor-level information, enabling effective floor-level global and room-level local optimizations as well as reliable loop closures across visually similar areas. We demonstrate a increase in accuracy and a reduction in computation time compared to the second-best method. Future work will integrate further semantic-relational concepts to enhance floor-based loop closure and optimization.

### v3

@@ VI-B Results and Discussions
| Method | Absolute Trajectory Error | Avg | |||
| E04 | E05 | E06 | E14 | ||
| HDL-SLAM [30] | 21.9 | 15.4 | 120.4 | 14.2 | 42.9 |
| ALOAM [5] | 28.2 | 22.5 | 130.1 | 24.9 | 51.4 |
| SCA-LOAM [31] | 28.6 | 23.1 | 130.4 | 25.3 | 51.9 |
| BALM [7] | 27.9 | 18.8 | 129.4 | 16.7 | 48.2 |
| S-Graphs+ [4] | 20.5 | 13.8 | 123.2 | 13.1 | 42.6 |
| Ours | 19.8 | 13.0 | 50.3 | 11.5 | 23.6 |
@@
| Dataset | Seq. Length [s] | Computation Time [ms] | |
| S-Graphs+ | Ours | ||
@@
Multi-Floor Dataset. Tab. III shows the Mean Map Entropy (MME) for the multi-floor dataset. Our method outperforms all the baselines by an average of when comparing the MME. Given the large area covered in these multi-floor experiments, all baselines fail at giving reasonable estimates, being unable to maintain the separation between the floor levels as well as accumulating significant drift. This can be assessed in Fig. 8, which shows the maps estimated by S-Graphs 2.0 and S-Graphs+ for LC1 (Top row (a-b) respectively), and in one of the multifloor Hilti datasets [29], bottom row (c-d). As these environments are highly aliased between floors, S-Graphs+’s loop closure gives a high number of false positives between levels. Our approach, given the floor segmentation, stairway detection, and floor-based loop closure, is able to maintain the accurate map while segmenting the keyframes and their connected layers based on the given floor-level. Fig. 9 shows qualitative results for CS2 demonstrating that our method is capable of segmenting the three floor-levels while maintaining a good map accuracy compared to S-Graphs+ (which again performs inaccurate loop closures between floor levels) and BALM. Tab. IV and Fig. 7 also validate the proper floor segmentation by comparing the IoU percentage of (height) values per floor-level from the maps generated by the algorithms versus the percentage ground truth time spend per floor-level. Observe that the IoU of our S-Graphs 2.0 correlates perfectly with the ground truth time, BALM presents a weaker correlation and the rest of baselines offer very poor results.
Remarkably, as reported in Tab. V, the average reduction of computation time in multi-floor dataset of S-Graphs 2.0 compared to S-Graphs+ is . As detailed before, this comes as a result of the hierarchical optimization and floor-level loop closure strategies of our S-Graphs 2.0.
@@ VII Conclusion
In this work, we present S-Graphs 2.0, a multi-layered hierarchical SLAM for large, multi-floor indoor environments. Our approach leverages semantic relationships and a robust floor detection module to assign floor-level information, enabling effective floor-level global and room-level local optimizations as well as reliable loop closures across visually similar areas. We demonstrate a increase in accuracy and a reduction in computation time compared to the second-best method. While our current floor detection relies on trajectory slope analysis and thus does not handle purely vertical transitions such as elevators, it effectively manages the more common case of stair-based floor transitions. Future work will integrate additional sensing modalities to handle elevator transitions, as well as further semantic-relational concepts to enhance floor-based loop closure and optimization.